# Modeling zero-inflated precipitation extremes

### v1

@@ 1 Introduction
Data characterized by excessive zeros and heavy-tailed distributions are prevalent in numerous fields, presenting unique challenges for statistical modeling. For instance, daily precipitation levels often exhibit a high frequency of zero values (days with no rainfall) along with a heavy-tailed distribution for positive rainfall amounts (Weglarczyk et al.,, 2005). Similarly, in the context of car insurance, the annual number of claims per client frequently includes a large proportion of policyholders with zero claims, whereas the remaining claims follow a skewed, heavy-tailed distribution (Christmann,, 2004). Another example can be found in healthcare, where the length of overnight hospital stay per patient often includes a significant clump-at-zero (patients discharged on the same day) and a right-skewed distribution for longer stays (Chen et al.,, 2007). In radio audience measurements, daily listening minutes for a given station are highly skewed and zero-inflated because of non-listeners (Couturier and Victoria-Feser,, 2010). These examples highlight the need for robust statistical models that are capable of handling datasets with excess zeros and heavy-tailed distributions.
@@
Zero inflation is a particularly common problem in daily and hourly precipitation data. For instance, rainfall measurements often include days with no precipitation, which results in a high frequency of zeros. In contrast, the positive values (rainy days) exhibited a heavy-tailed distribution owing to occasional extreme rainfall events. In various studies, researchers have attempted to model only the extreme right tail using the peak over threshold (POT) method, which models exceedances beyond a high threshold. When the threshold is sufficiently high, the distribution of exceedances often approximates the Generalized Pareto Distribution (GPD). A significant limitation of GPD is that it can only model observations that exceed a specified high threshold, leaving uncertainty about the behavior of the remaining observations below this threshold. Various attempts have been made to model the entire range of data using a mixture of models. We refer to Frigessi et al., (2002), who proposed a dynamic mixture model for unsupervised tail estimation. Using a weight function, the model combined the light- and heavy-tailed densities of the GPD. Carreau and Bengio, (2009) proposed a hybrid Pareto distribution by extending GPD that can be used in a mixture model to model the entire real axis. Scarrott and MacDonald, (2012) offers a comprehensive overview of mixture modeling and its applications by introducing the mixture form of two models below and above the specified threshold. Papastathopoulos and Tawn, (2013) introduces a new extension of GPD that models low-threshold exceedances and retains the same tail behavior as GPD. A drawback of these approaches is the need to select an appropriate threshold and an increased number of parameters, which complicates the inference. Following Papastathopoulos and Tawn, (2013), Naveau et al., (2016) proposed extensions of GPD that are sufficiently flexible to model the entire range of positive rainfall intensities without a threshold selection step. They employed censored likelihood methods to handle low precipitation values effectively. Stein, (2021) introduced both tails of the flexible distribution to model the tails of the data distribution by the tail index of GDP. The issue of zero inflation in precipitation data has been largely overlooked despite its critical impact on statistical modeling and climate analyses.
@@
To address the challenge of modeling zero inflation and heavy tails together, this study proposes a Zero-inflated Extended Generalized Pareto Distribution (ZIEGPD) for modeling zero-inflated precipitation data. ZIEGPD has a tail index parameter similar to that of GPD. Therefore, the proposed model is well suited for modeling heavy-tailed data and excessive zeros, and the distribution bypasses the threshold choice step, making it an ideal candidate for describing dry and wet events simultaneously. By extending the GPD to a zero-inflated framework, our model can simultaneously estimate the probability of zero precipitation and the distribution of positive rainfall amounts, providing a comprehensive approach for analyzing zero-inflated heavy-tailed data. Several constructions have been proposed in the literature to handle zero inflation in data. The most commonly applied method for handling zero inflation was introduced by Lambert, (1992) by mixing a two-part model specifically for continuous data zero-inflated models to account for excess zeros. Ahmad et al., (2024), Ahmad and Arshad, (2024) proposed a flexible model for zero-inflated heavy-tailed discrete data. Section  2 describes the proposed model as a natural extension of the aforementioned models for zero-inflated heavy-tailed continuous data. The proposed ZIEGPD is quite flexible at the bulk part of the distribution, while correctly modeling the zero proportion and tail of the distribution. For estimation purposes, we introduce maximum likelihood and Bayesian estimation paradigms.
@@
The remainder of this paper is organized as follows. Section 2 provides the background of zero-inflated mixture models and introduces the proposed model. Section 3 discusses the inferential procedures. Detailed simulation results are provided in Section  4. The real-data application findings are explained in Section  5. Section 6 concludes the paper.
@@ 2 Modeling framework
where $(a)_{+}=\max(a,0)$, $\sigma>0$ and $-\infty<\xi<+\infty$ represent the scale and shape parameters of the distribution, respectively.
@@
For the development of the models proposed in this study, we following the way outlined by Naveau et al., (2016), which uses the idea of integral transformation to simulate random draws from a GPD. Specifically, their method involves generating GPD random variables by applying the inverse of the GPD CDF in (1), denoted by $\mathcal{H}_{\sigma,\xi}^{-1}(U)$, where $U\sim\mathcal{U}(0,1)$ is a random variable uniformly distributed on the interval $(0,1)$. We can construct a family of models through the idea of inverse transformation for random variables
@@
[…]
@@
The CDF $\mathcal{W}(\mathcal{H}(z;\sigma,\xi))$, referred to as the CDF of Extended GPD (EGPD), is valid only for modeling non-zero observations or low threshold exceedances. For data containing excessive zeros, a censored likelihood method is required along with rounding zero values to a low non-zero number (for instance, round to 0.1 in hourly precipitation data) (Naveau et al.,, 2016).
In numerous practical scenarios, datasets frequently exhibit an excessive number of zeros. These excessive zeros complicate precise statistical analysis, as standard statistical models are typically not designed to handle such cases. Thus, it is crucial to explore the underlying causes of these zeros. This problem can be tackled through a mixture of zero inflation with non-zero observations by a mixture model.
@@ 2.2 Zero inflated extended generalized Pareto (ZIEGP) distribution
By integrating the density function of ZIEGP distribution given in (7), we found the CDF of ZIEGP distribution as
@@
where $z\geq 0$, $0\leq\pi\leq 1$.
$q_{p^{*}}=H^{-1}(p^{*})\left\{\begin{array}[]{ll}\frac{\sigma}{\xi}\left\{\left(1-\mathcal{W}^{-1}(p^{*})\right)^{-\xi}-1\right\},&\mbox{if }\xi>0\\ \\ -\sigma\log\left(1-\mathcal{W}^{-1}(p^{*})\right),&\mbox{if }\xi=0,\\ \end{array}\right.$ (9)
@@
The expression (7) with incorporation of $\mathcal{W}(u;\Psi)=u^{\kappa}$, $\Psi=\kappa>0$ leads to the density function of ZIEGPD with parameters $(\pi,\kappa,\sigma,$ and $\xi)$, where $\pi$ is the parameter for proportion of zeros, $\kappa$ is the shape parameter for the bulk part of the distribution and $\sigma$ and $\xi$ are usual scale and shape parameters. Figure 1 (top-left) illustrates the density behavoir with fixed values for $(\pi=0.60,\sigma=1$ and $\xi=0.5)$ while varying the bulk part behavior by changing $\kappa$ to 1, 2, and 5. When $\kappa$ is set to 1, the model reduces to the EGPD with addition of zero inflation. Increasing $\kappa$ introduces greater flexibility in the lower tail behavior along with same amount of zero inflation and without compromising the characteristics of the upper tail, as evidenced in Figure 1 (top-left) when $\kappa$ is set to 5.
The expression (7) with combination of $\mathcal{W}(u;\Psi)=1-B_{\delta}\{(1-u)^{\delta}\}$, $\Psi=\delta>0$ leads to another density function of ZIEGPD with parameters $(\pi,\delta,\sigma,$ and $\xi)$, where $\delta$ is characterizes the central portion of the distribution and enhancing flexibility in that region. In contrast, other parameters have the same role as we discuss above. With varying $\delta$ values, Figure 1 (top-right) shows different behaviors. For large values of $\delta$, the distribution behaves similarly to GPD with an additional amount of zeros at the origin. Decreasing $\delta$ values only shows flexibility at the central part of the distribution without affecting the upper tail behavior and keeping the same zero inflation at the origin.
@@
To overcome the drawback seen in $\mathcal{W}(u;\Psi)=1-B_{\delta}\{(1-u)^{\delta}\}$, $\Psi=\delta>0$, we introduce flexibility at bulk part of distribution by additional parameter $\kappa$, the expression (7) with combination of $\mathcal{W}(u;\Psi)=[1-B_{\delta}\{(1-u)^{\delta}\}]^{\kappa/2}$, $\Psi=(\delta,\kappa)$ with $\delta>0$ and $\kappa>0$ leads to another flexible density function of ZIEGPD with parameters $(\pi,\delta,\kappa,\sigma,$ and $\xi)$, where $\delta$ is characterizes the central part of the distribution and $\kappa$ creates flexibility at bulk part of the distribution. With varying $\kappa$ values, Figure 1 (bottom) shows flexibility at the lower part of the distribution, even when varying $\delta$ to high values. For instance, at $(\delta=5,\kappa=10)$, the clear flexibility at the lower part can be observed with the inflation of zeros in the model.
@@ 3 Inference for ZIEGP model
This section introduces the methods for estimating the ZIEGP model. For the perspective of obtaining better results even in case of nonstationary data.
@@
After solving the derivative involve in (10) for each $\mathcal{W}$ choice, the expressions for the density functions are provide in Appendix Appendix. The log likelihood is defined as
@@
In case of non-stationary data, we also developed the non- stationary version of proposed models. It is a semi-parametric regression model that assumes a parametric distribution for the response variable, with the option for the distribution parameters to change according to explanatory variables or random effects. In non-stationary framework, the observation $z_{t},t=1,\cdots,n$ and fitted to distribution $h(z_{t}|\theta^{t})$ with $\theta^{t}=(\theta_{1t},\cdots,\theta_{pt})$ where $p$ is the number of distribution parameters at time $t$. To relate the parameters of the distribution to time $t$ or vector of covariates $(\boldsymbol{x_{i}})$ of dimension $p$, we consider a generic structure of additive predictors of the form
$\eta_{j}^{\theta^{t}}=\beta_{j0}^{{\theta^{t}}}+f_{j1}^{{\theta^{t}}}(\boldsymbol{x}_{ij1})+\dots+f_{jk}^{{\theta^{t}}}(\boldsymbol{x}_{ijk})+\dots+f_{jK_{j}}^{{\theta^{t}}}(\boldsymbol{x}_{ijK_{j}}),$ (12)
Our proposed models do not incorporate covariates at this stage; however, the developed methodology remains adaptable for their inclusion if desired. For the classical approach, we implement the proposed models using our own developed R code and incorporate it as a custom families in the evgam (Youngman,, 2022) R package. For Bayesian estimation, we introduce the models as an add-on in the bamlss R package (Umlauf et al.,, 2021). For all parametric, linear effects, we follow Klein et al., (2015) and assume a flat, noninformative prior for intercept, representing a limiting case of a multivariate Gaussian prior with high dispersion. This approach can provide regularization for high-dimensional parameter vectors.
@@ 4.1 Model based simulation
The boxplots of estimated parameters of all three models through MLE and Bayesian are shown in Figure 2, for sample size $n=500$ in (left panel) and sample size $n=1000$ in (right panel). The top panel (left-to-right) illustrates the results obtained by fitting Model $M_{1}$ using both classical and Bayesian paradigms. Boxplots depict the estimation performance of the true parameters across all configurations. Both methods accurately estimate the true parameters, demonstrating robustness across scenarios. Furthermore, the proportion of zeros and bulk parameter $\kappa$ are consistently estimated even in cases where the values $\xi$ are relatively large, i.e., $\xi=0.4$.
@@
((d))
((e))
((f))
Figure 2: Box plots showing the maximum likelihood estimates and Bayesian estimates of model parameters for each model type, based on sample sizes of $n=500$ and $n=1000$, with $10^{4}$ replications under varying parameter settings. Panels (a), (c), and (e) represent model types $M_{1}$, $M_{2}$, and $M_{3}$ for $n=500$, respectively, while panels (b), (d), and (f) correspond to model types $M_{1}$, $M_{2}$, and $M_{3}$ for $n=1000$.
The middle and bottom panels (left-to-right) present boxplots of the estimated parameters obtained by fitting models $M_{2}$ and $M_{3}$, respectively. In both models, the parameter $\delta$, which governs the skewness, proves challenging to estimate accurately. The estimates of $\delta$ exhibit greater variability across configurations, regardless of whether the classical or Bayesian method is employed. This increased variation highlights the inherent difficulty in reliably estimating parameters associated with skewness in these models. Such estimation difficulties of skewness parameter of models were also noticed in literature by Sartori, (2006), Ribereau et al., (2016). In the estimation of remaining parameters, the MLE performs better in small samples, while Bayesian estimates look more stable in large sample settings. Again, the parameter for the zeros proportion is estimated correctly with both methods regardless of the sample size choice.

### v2

@@ 1 Introduction
Data characterized by excessive zeros and heavy-tailed distributions are prevalent in numerous fields, presenting unique challenges for statistical modeling. For instance, daily precipitation levels often exhibit a high frequency of zero values (days with no rainfall) along with a heavy-tailed distribution for positive rainfall amounts (Weglarczyk et al.,, 2005). Similarly, in the context of car insurance, the annual number of claims per client frequently includes a large proportion of policyholders with zero claims, whereas the remaining claims follow a skewed, heavy-tailed distribution (Christmann,, 2004). Another example can be found in healthcare, where the length of an overnight hospital stay per patient often includes a significant clump at zero (patients discharged on the same day) and a right-skewed distribution for longer stays (Chen et al.,, 2007). In radio audience measurements, daily listening minutes for a given station are highly skewed and zero-inflated because of non-listeners (Couturier and Victoria-Feser,, 2010). These examples underscore the need for robust statistical models that can effectively handle datasets with excess zeros and heavy-tailed distributions.
@@
Zero-inflation is a particularly common problem in daily and hourly precipitation data. For instance, rainfall measurements often include days with no precipitation, which results in a high frequency of zeros. In contrast, the positive values (rainy days) exhibited a heavy-tailed distribution owing to occasional extreme rainfall events. In various studies, researchers have attempted to model only the extreme right tail using the peak over threshold (POT) method, which models exceedances beyond a high threshold. When the threshold is sufficiently high, the distribution of exceedances often approximates the Generalized Pareto Distribution (GPD). A significant limitation of GPD is that it can only model observations that exceed a specified high threshold, leaving uncertainty about the behavior of the remaining observations below this threshold. Various attempts have been made to model the entire range of data using a mixture of models. Kedem et al., (1990) highlighted that rainfall is best represented by a mixed distribution, with a discrete spike at zero for dry periods and a continuous distribution (lognormal or gamma) for wet periods. Frigessi et al., (2002) proposed a dynamic mixture model for unsupervised tail estimation. Using a weight function, the model combined the light- and heavy-tailed densities of the GPD. Carreau and Bengio, (2009) proposed a hybrid Pareto distribution by extending GPD that can be used in a mixture model to model the entire real axis. Scarrott and MacDonald, (2012) offers a comprehensive overview of mixture modeling and its applications by introducing the mixture form of two models below and above the specified threshold. Papastathopoulos and Tawn, (2013) introduces a new extension of GPD that models low-threshold exceedances and retains the same tail behavior as GPD. A drawback of these approaches is the need to select an appropriate threshold and an increased number of parameters, which complicates the inference. Following Papastathopoulos and Tawn, (2013), Naveau et al., (2016) proposed extensions of GPD that are sufficiently flexible to model the entire range of positive rainfall intensities without a threshold selection step. They employed censored likelihood methods to handle low precipitation values effectively. Stein, (2021) introduced both tails of the flexible distribution to model the tails of the data distribution by the tail index of GDP. Wilson et al., (2022) proposed a deep extreme mixture model that offers a unified framework for better capturing moderate events, including zero and extreme events, while allowing flexible thresholding for spatio-temporal forecasting of precipitation. Haruna et al., (2025) applied a non-stationary extended GPD to model long-term trends in both the bulk and the extremes of daily precipitation distributions in Switzerland, while dry events were treated separately in their modeling strategy. Overall, the issue of zero inflation in precipitation data has been largely overlooked despite its critical impact on statistical modeling.
@@
To address the challenge of modeling zero inflation and heavy tails together, this study proposes a Zero-inflated Extended Generalized Pareto Distribution (ZIEGPD) for modeling zero-inflated precipitation data. ZIEGPD has a tail index parameter similar to that of GPD. Therefore, the proposed model is well-suited for modeling heavy-tailed data and excessive zeros, and the distribution bypasses the threshold choice step, making it an ideal candidate for describing dry and wet events simultaneously. By extending the EGPD to a zero-inflated framework, our model can simultaneously estimate the probability of zero precipitation and the distribution of positive rainfall amounts, providing a comprehensive approach for analyzing zero-inflated heavy-tailed data. Several constructions have been proposed in the literature to handle zero inflation in data. The most commonly applied method for handling zero inflation was introduced by Lambert, (1992) by mixing a two-part model specifically for continuous data zero-inflated models to account for excess zeros. Ahmad et al., (2024), Ahmad and Arshad, (2024) proposed a flexible model for zero-inflated heavy-tailed discrete data.Dzupire et al., (2018)showed that modeling rainfall occurrence and intensity separately tends to underestimate the variability inherent in rainfall data, and they proposed a Poisson–Gamma model that jointly accounts for both zero and non-zero rainfall. Ensemble precipitation forecasts are often unreliable due to zero-inflation, heteroscedasticity, and rare extreme events. To address limitations of separate modeling of occurrence and intensity, Scheuerer and Hamill, (2015) proposed a censored, shifted gamma distribution (CSGD) model that jointly models zeros and precipitation amounts, improving probabilistic forecast.
Section 2 describes the proposed model as a natural extension of the aforementioned models for zero-inflated heavy-tailed continuous data. The proposed ZIEGPD is quite flexible at the lower tail of the distribution, while correctly modeling the zero proportion and tail of the distribution. For estimation purposes, we introduce maximum likelihood and Bayesian estimation paradigms.
@@
The remainder of this paper is organized as follows. Section 2 provides the background of zero-inflated mixture models and introduces the proposed model. Section 3 discusses the inferential procedures. Detailed simulation results are provided in Section 4. The real-data application findings are explained in Section 5. Section 6 concludes the paper.
@@ 2 Modeling framework
where $(a)_{+}=\max(a,0)$, $\sigma>0$ and $\infty<\xi<+\infty$ represent the scale and shape parameters of the distribution, respectively.
@@
For the development of the models proposed in this study, we follow the way outlined by Naveau et al., (2016), which uses the idea of integral transformation to simulate random draws from a GPD. Specifically, their method involves generating GPD random variables by applying the inverse of the GPD CDF in (1), denoted by $\mathcal{H}_{\sigma,\xi}^{-1}(U)$, where $U\sim\mathcal{U}(0,1)$ is a random variable uniformly distributed on the interval $(0,1)$. We can construct a family of models through the idea of inverse transformation for random variables
@@
$\lim_{u\to 0}\frac{\bar{\mathcal{W}}(1-u)}{u}$ $=$ $a,\quad\text{for some constant }a>0\text{ (upper tail behavior)},$ (3)
$\lim_{u\to 0}\frac{\mathcal{W}(u)}{u^{\kappa}}$ $=$ $c,\quad\text{for some constant }c>0\text{ (lower tail behavior)}.$ (4)
For CDF of $\mathcal{W}$, we consider three families proposed by Naveau et al., (2016), denoted by $\mathcal{W}(\cdot,\Psi)$,that satisfy the conditions outlined in (3) and (4)
@@
The CDF $\mathcal{W}(\mathcal{H}(z;\sigma,\xi))$, referred to as the CDF of Extended GPD (EGPD), is valid only for modeling non-zero observations or low threshold exceedances. For nonzero precipitation data, a censored likelihood approach is recommended to improve estimation, with low nonzero values rounded (e.g., to 0.1 for hourly data) to reduce discretization effects (Naveau et al.,, 2016).
In numerous practical scenarios, datasets frequently exhibit an excessive number of zeros. These excessive zeros complicate precise statistical analysis, as standard statistical models are typically not designed to handle such cases. Thus, it is essential to employ modeling approaches that reduce bias and properly capture the data structure. One effective approach is to use mixture models that jointly account for zero inflation and non-zero observations.
@@ 2.2 Zero inflated extended generalized Pareto (ZIEGP) distribution
By integrating the density function of ZIEGPD given in (7), we found the CDF of ZIEGPD as
@@
where $z\geq 0$, $0\leq\pi\leq 1$. The quantile function is defined as
[…]
@@
The expression (7) with incorporation of $\mathcal{W}(u;\Psi)=u^{\kappa}$, $\Psi=\kappa>0$ leads to the density function of ZIEGPD with parameters $(\pi,\kappa,\sigma,$ and $\xi)$, where $\pi$ is the parameter for proportion of zeros, $\kappa$ is the shape parameter for the lower tail of the distribution and $\sigma$ and $\xi$ are usual scale and shape parameters. Figure 1 (top-left) illustrates the density behavior with fixed values for $(\pi=0.60,\sigma=1$ and $\xi=0.5)$ while varying the lower tail behavior by changing $\kappa$ to 1, 2, and 5. When $\kappa$ is set to 1, the model reduces to the GPD with the addition of zero inflation. Increasing $\kappa$ introduces greater flexibility in the lower tail behavior along with same amount of zero inflation and without compromising the characteristics of the upper tail, as evidenced in Figure 1 (top-left) when $\kappa$ is set to 5.
The expression (7) with combination of $\mathcal{W}(u;\Psi)=1-B_{\delta}\{(1-u)^{\delta}\}$, $\Psi=\delta>0$ leads to another density function of ZIEGPD with parameters $(\pi,\delta,\sigma,$ and $\xi)$, where $\delta$ is characterizes the central portion of the distribution and enhancing flexibility in that region. In contrast, other parameters have the same role as we discussed above. With varying $\delta$ values, Figure 1 (top-right) shows different behaviors. For large values of $\delta$, the distribution behaves similarly to GPD with an additional number of zeros at the origin. Decreasing $\delta$ values only show flexibility at the central part of the distribution without affecting the upper tail behavior and keeping the same zero inflation at the origin, while the lower tail remains difficult to estimate reliably from the data (see Figure 1 (top-right)). This drawback has also been noted by Naveau et al., (2016).
@@
To overcome the drawback seen in $\mathcal{W}(u;\Psi)=1-B_{\delta}\{(1-u)^{\delta}\}$, $\Psi=\delta>0$, we introduce flexibility at lower tail of distribution by additional parameter $\kappa$, the expression (7) with combination of $\mathcal{W}(u;\Psi)=[1-B_{\delta}\{(1-u)^{\delta}\}]^{\kappa/2}$, $\Psi=(\delta,\kappa)$ with $\delta>0$ and $\kappa>0$ leads to another flexible density function of ZIEGPD with parameters $(\pi,\delta,\kappa,\sigma,$ and $\xi)$, where $\delta$ is characterizes the central part of the distribution and $\kappa$ creates flexibility at lower tail of the distribution. With varying $\kappa$ values, Figure 1 (bottom) shows flexibility at the lower part of the distribution, even when varying $\delta$ to high values. For instance, at $(\delta=5,\kappa=10)$, the clear flexibility at the lower part can be observed with the inflation of zeros in the model.
@@ 3 Inference for ZIEGP model
This section introduces the methods for estimating the ZIEGP model. For the purpose of obtaining better results, even in the case of nonstationary data.
@@
After solving the derivative involved in (10) for each $\mathcal{W}$ choice, the expressions for the density functions are provided in Appendix A. The log likelihood is defined as
@@
In case of non-stationary data, we also developed the non-stationary version of the proposed models along with classical and Bayesian paradigms, see Appendix B. From an application perspective, the proposed models do not consider incorporating covariates; however, the developed methodology remains adaptable for their inclusion if desired. For the classical approach, we implement the proposed models using our own developed R code and incorporate it as a custom families in the evgam (Youngman,, 2022) R package. For Bayesian estimation, we introduce the models as an add-on in the bamlss R package (Umlauf et al.,, 2021). For all parametric, linear effects, we follow Klein et al., (2015) and assume a flat, noninformative prior for the intercept, representing a limiting case of a multivariate Gaussian prior with high dispersion. This approach can provide regularization for high-dimensional parameter vectors.
@@ 4.1 Model based simulation
The boxplots of estimated parameters of all three models through MLE and Bayesian are shown in Figure 2, for sample size $n=1000$. Figure S.1 in the supplement file reports the results for $n=500$. The top-left panel in both figures illustrates the results obtained by fitting Model $M_{1}$ using both classical and Bayesian paradigms. Boxplots depict the estimation performance of the true parameters across all configurations. Both methods accurately estimate the true parameters, demonstrating robustness across scenarios. Furthermore, the proportion of zeros and bulk parameter $\kappa$ are consistently estimated even in cases where the values $\xi$ are relatively large, i.e., $\xi=0.4$.
@@
Figure 2: Box plots showing the maximum likelihood estimates and Bayesian estimates of model parameters for each model type, based on sample sizes of $n=1000$, with $10^{4}$ replications under varying parameter settings.
The right and bottom panels in both figures present boxplots of the estimated parameters obtained by fitting models $M_{2}$ and $M_{3}$, respectively. In both models, the parameter $\delta$, which governs the skewness, proves challenging to estimate accurately. The estimates of $\delta$ exhibit greater variability across configurations, regardless of whether the classical or Bayesian method is employed. This increased variation underscores the inherent difficulty in reliably estimating parameters related to skewness in these models. Such estimation difficulties of the skewness parameter of models were also noticed in literature by Sartori, (2006), Ribereau et al., (2016). In the estimation of remaining parameters, the MLE performs better in small samples, while Bayesian estimates look more stable in large sample settings. Again, the parameter for the zero proportion is estimated correctly with both methods, regardless of the sample size choice.